OpenStudy (anonymous):

@TheSmartOne The first five terms of a sequence are shown below: 6, 9, 12, 15, 18 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 6n - 3; n ≥ 1 f(n) = 6n + 3; n ≥ 1 f(n) = 3n + 3; n ≥ 1 f(n) = 3n - 3; n ≥ 1

OpenStudy (anonymous):

b?

TheSmartOne (thesmartone):

Nope

OpenStudy (ddcamp):

What's the difference between each term?

TheSmartOne (thesmartone):

@Supermariofan153 Please no direct answers. Its against the CoC http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

each is 3 more than each other

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@DDCamp

OpenStudy (ddcamp):

The difference between each term will be what you multiply n by. (So if n goes up by one, the term goes up by 3). Does that make sense?

OpenStudy (anonymous):

yes @DDCamp

TheSmartOne (thesmartone):

Since 6 is the first number then when n=1 then f(n)=6

TheSmartOne (thesmartone):

So when n=1, f(n) = 6 n=2, f(n)= 9 and so on

TheSmartOne (thesmartone):

So basically plug in those numbers and you got it :)

OpenStudy (anonymous):

ok so basically c @TheSmartOne

TheSmartOne (thesmartone):

Yup :D

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 @heeeelllooo

OpenStudy (anonymous):

thank you @TheSmartOne

TheSmartOne (thesmartone):

Np :)

OpenStudy (anonymous):

A sequence has its first term equal to 3, and each term of the sequence is obtained by adding 5 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence? f(1) = 3 and f(n) = f(n - 1) + 5; n > 1 f(1) = 5 and f(n) = f(n - 1) + 3; n > 1 f(1) = 3 and f(n) = f(n - 1) + 5n; n > 1 f(1) = 5 and f(n) = f(n - 1) + 3n; n > 1

OpenStudy (anonymous):

@TheSmartOne c o. d